# MVGT: A Multi-view Graph Transformer Based on Spatial Relations for EEG Emotion Recognition

### v4

@@ Abstract
Index Terms:
EEG, emotion recognition, graph transformer, spatial encoding
I Introduction
Figure 1: Overall structure of MVGT. (a) represents the process of brain region encoding, centrality encoding and geometric structure encoding. (b) depicts the process of calculating inter-channel correlations based on the attention mechanism and geometric structure encoding. “Recycling” refers to the iterative refinement (see II-D).
@@ I Introduction
Affective computing is commonly employed for the analysis of emotional states through Human-Computer Interaction (HCI) systems, which collect multimodal data from subjects, including voice signals, self-reports, body gestures and physiological signals. Among these modalities, physiological signals offer distinct advantages as they are directly derived from the subjects’ mental states, making it difficult for them to disguise or conceal their emotions. Scalp electroencephalography (EEG), a physiological signal usually used to analyze the cognitive functions of the human brain, is collected using noninvasive electrodes on the scalp [1]. Nowadays, due to its high temporal resolution, portability, and affordability, this method is widely employed to study brain changes in response to emotional stimuli [2].
@@
EEG signals contain rich emotional information in the temporal, frequency, and spatial domains. How to reasonably and effectively utilize information from each domain is a major challenge. Therefore, recent studies have attempted to combine multi-domain information to capture complex EEG features [3, 4, 5]. Although these works have achieved good performance, their limited use of domain-specific information has constrained model performance.
@@
The frequency domain features of EEG commonly include power spectral density (PSD), differential entropy (DE), differential asymmetry (DASM), and others. Among these, DE has been proven to be the simplest and most effective in EEG emotion recognition tasks [6]. DE has the ability to distinguish EEG patterns between high-frequency and low-frequency energy [7].
@@
Given the high temporal resolution of EEG signals, many studies use temporal models to extract sequential information. Recurrent neural networks (RNNs) have significantly improved emotion recognition performance by modeling temporal dependencies [8, 9]. However, their serial structure poses challenges for parallelization. Convolutional neural networks (CNNs) have also been applied to extract temporal features, but their performance is restricted by limited receptive fields [10, 3]. To overcome these difficulties, [4] employs the attention mechanism to capture relationships between time points; however, this method may suffer performance degradation due to time-unaligned events at a single moment.
@@
The spatial domain of EEG is equally crucial for understanding emotional states, as emotions often involve distributed neural circuits rather than isolated regions [11]. Asymmetric activity between the left and right hemispheres can reflect changes in valence and arousal [12]. Message-passing-based graph neural network (MPGNN) methods can effectively capture structural relationships between EEG channels [13, 14, 10, 3]. However, using MPGNN may pose potential risks of over-smoothing and over-squashing, which could hinder its ability to learn complex representations [15]. Additionally, these methods often underutilize the geometric and anatomical structural relationships inherent in EEG spatial information.
@@
To address these limitations and fully exploit features in the temporal, frequency, and spatial domains, we propose a multi-view graph transformer (MVGT) based on spatial relations for EEG emotion recognition. Graph transformers offer robust and flexible spatial perception while mitigating issues like over-smoothing and over-squashing [15]. For the frequency domain, we use the embedding of DE as the feature representation. For the temporal domain, inspired by [16], we expand the temporal receptive field by embedding entire continuous time segments as tokens and employ feed-forward neural networks to capture temporal patterns. In the spatial domain, we incorporate several simple yet effective spatial encodings, brain region encoding, centrality encoding and geometric structure encoding, into the model to help it capture effective spatial structures. These encodings allow MVGT to adaptively model correlations between EEG channels, delivering improved performance and deeper insights into emotional states.
II Methods
II-A Problem Definition
@@
II-B Temporal Embedding
Temporal information is critical in time-series processing. [4] treated the features of different channels at the same time points as tokens and employed the attention mechanism to extract temporal correlations between them. However, due to anisotropic volume conduction characteristics [17] in human brain tissues, there may be different physical meanings between channels at the same time step, resulting time-unaligned events that subsequently degrade performance. Other approaches [3, 10] utilized CNNs to extract temporal information along the time axis from continuous EEG segments, with the receptive field limited by the size of the convolution kernel. Inspired by [16], we broaden the receptive field by embedding entire time segments as tokens rather than single time points. EEG signals are segmented using overlapping sliding windows of size $T$, and these segments are treated as tokens. After processing with sliding windows, we obtain $\tilde{\bm{X}}=\left(\tilde{\bm{X}}_{1},\tilde{\bm{X}}_{2},\cdots,\tilde{\bm{X}}_{S}\right)\in\mathbb{R}^{S\times n\times Tf}$, where $S$ denotes the number of continuous EEG segments, $n$ is the number of channels, and $f$ denotes the dimension of DE features. Refer to the Appendix B for more temporal embedding details.
According to the universal approximation theorem [18], the feed-forward neural network (FFN), as the basic module of the Transformer [19] encoder, can learn the intrinsic properties to describe a time series. By processing continuous time segments as input, the FFN extracts temporal features independently for each channel, providing more advantageous predictive representations compared to self-attention [16].
@@ II-B Temporal Embedding
II-C Spatial Encoding
@@
II-C1 Brain Region Encoding
Neuroscience research indicates that specific brain regions activate together during high-level cognitive processes [20]. In EEG emotion recognition, incorporating relevant neuroscience findings can typically enhance recognition accuracy. For example, studies like [13] and [21] utilized the asymmetry of neural activity between the left and right hemispheres to enhance recognition performance. [10] divided EEG channels into different regions and combined intra-region convolution with inter-region convolution to capture spatial relationships. Building on [10], we propose four learnable BRE schemes based on neuroscience research, ranging from coarse-grained to fine-grained encodings, namely: LOBE, FRONTAL, GENERAL, and HEMISPHERE. The four schemes mentioned above are shown in Fig. 2.
@@ II-C1 Brain Region Encoding
II-C2 Structure Encoding
EEG channels are distributed in the 3D space, and the functional connectivity between them remains imprecisely defined. Therefore, inspired by [22], we model the relationships between EEG channels as a fully connected directed graph to avoid making specific assumptions about functional connectivity. Firstly, let $\phi(i,j)$ represent the Euclidean distance between node $i$ and node $j$, and encode $\phi(i,j)$ using a set of Gaussian basis functions [22, 23]. Let $\bm{b}_{k}\in\mathbb{R}^{n\times n}$ denote one of the Gaussian basis functions. The element $(i,j)$ of this function can be expressed as:
@@ II-C2 Structure Encoding
[…]
@@
Unlike previous symmetric adjacency matrices [4, 13], our directed spatial encoding matrix allows the model to learn distinct correlations for $\left(i,j\right)$ and $\left(j,i\right)$. Letting $l$ denote the model depth, $s$ denote the sample index and $m$ denote the index of multi-head attention, the multi-head attention module (MHA) can be represented as:
@@
II-D Implementation Details of MVGT
In this section, we describe the overall architecture of the model, as illustrated in Fig. 1. For better optimization, we first apply GraphNorm [24] to normalize the input features to a range between 0 and 1. The encodings could be characterized as below:
@@ II-D Implementation Details of MVGT
We employ a Pre-LN structure, applying layer normalization (LN) before the MHA and the FFN. This design choice is supported by prior research [25], which demonstrates that Pre-LN structures produce more stable gradients, facilitating faster and more reliable convergence. To further prevent overfitting, dropout is employed during training. The process is expressed mathematically as:
@@
Inspired by [22, 26], we apply iterative refinement by recursively feeding the outputs of the model back into the same modules, denoted as “recycling” in Fig. 1. The iterative approach refines the model’s ability to discriminate encoded information and understand emotional patterns, thereby helping the model capture more effective details.
III Experiments
III-A Datasets and Settings
For our experiments, we select the SEED [7] and SEED-IV [27] datasets to evaluate the effectiveness of our model. These datasets consist of EEG signals recorded from subjects watching emotion-eliciting videos. Following the settings of previous studies [7, 21, 14, 13, 3, 4, 27, 5], we use pre-computed differential entropy (DE) features for the recognition task. Time segments are extracted using an overlapping sliding window of size $T$, consistent with [3], with $T$ set to 5. For the SEED dataset, we use the first 9 trials of each subject as the training set and the last 6 trials as the test set, as done in previous research. The DE features are computed using five frequency bands extracted from 1s nonoverlapping windows. The model performance is evaluated based on the average accuracy and standard deviation across all subjects over two sessions of EEG data. Similarly, for the SEED-IV dataset, we allocate the first 16 trials to training and the last 8 trials to testing. The DE features for SEED-IV are calculated using 4s windows. The performance of our model is assessed using data from all three sessions. Refer to Appendix A for DE feature extraction.
Table I: The classification accuracies (mean/std) on SEED and SEED-IV. MVGT-L, MVGT-G, MVGT-H, MVGT-F: MVGT using LOBE, GENERAL, HEMISPHERE and FRONTAL scheme. The best results are highlighted in bold, and the second-best results are underlined.
@@ III-A Datasets and Settings
DGCNN [14] 90.40/08.49 69.88/16.29
BiDANN [9] 92.38/07.04 70.29/12.63
BiHDM [21] 93.12/06.06 74.35/14.09
R2G-STNN [8] 93.34/05.96 -
RGNN [13] 94.24/05.95 79.37/10.54
MD-AGCN [3] 94.81/04.52 87.63/05.77
EmoGT [4] 95.02/05.99 91.20/09.60
MV-SSTMA [5] 95.32/3.05 92.82/5.03
@@
During experiments, the hidden dimension is set to 64 and the number of Gaussian basis functions is 32. The number of MHA layers is 4 and the number of attention heads is 2. We employ iterative refinement by recursively feeding the output back into the model for three times. We set the batch size to 32 and the learning rate within the range of 3e-5 to 3e-3. Cross-entropy is used as the loss function, and AdamW [28] is employed as the optimizer with a weight decay rate of 0.1.
@@
III-B Results Analysis
We compare the classification results on the SEED and SEED-IV datasets with recent state-of-the-art models, as shown in Table I. Under identical experimental conditions, our proposed model demonstrated significantly better results than the baselines. On the SEED dataset, the FRONTAL scheme achieved the highest classification accuracy of 96.55%, representing a 1.23% improvement over the best baseline model. Similarly, on the SEED-IV dataset, the GENERAL scheme reached an accuracy of 94.03%, exceeding the best baseline by 1.21%. The differences in accuracy across region division schemes likely result from their data-dependent nature. This suggests that designing data-specific division schemes can enable the model to better utilize data characteristics, thereby enhancing recognition performance. Conversely, improper schemes may result in poor performance.
Fig. 3 shows the confusion matrices for MVGT-F on SEED and MVGT-G on SEED-IV. For SEED, the model achieves 97.70% accuracy for positive emotions but performs less effectively on negative emotions (95.61%). Misclassification rates are low, with 0.76% of positive samples misclassified as negative and 1.08% of negative samples as positive, demonstrating effectiveness in distinguishing valence changes. For SEED-IV, the model achieves its best performance on neutral emotions (95.77%) but slightly lower accuracy on happy emotions (90.30%), likely due to the GENERAL scheme’s sensitivity to balanced emotions.
III-C Ablation Study
To validate the effectiveness of our proposed method, we conduct ablation experiments on MVGT-F (based on SEED) and MVGT-G (based on SEED-IV). These experiments incrementally remove four key components: centrality encoding (CE), brain region encoding (BRE), geometric structure encoding (GSE), and the inverted temporal embedding (Inverted). GSE is incorporated into the attention mechanism as a bias (as detailed in (8)).
Table II: Ablation study for the classification accuracies (mean/std) of MVGT-F on SEED and MVGT-G on SEED-IV. Symbol “✓” indicates the component is employed.
@@ III-C Ablation Study
The term “Inverted” refers to embedding continuous time segments as tokens (described in Section II-B), whereas the default approach treats multi-channel information from a single time point as tokens.
@@
III-D Visualization of Inter-channel Relations
@@
On the SEED dataset, the parameters suggest that emotional patterns involve activity across multiple brain regions. The parameters from head-1 in MVGT-F suggest that brain activity is primarily concentrated in the temporal and frontal regions on the lateral side of the brain, which aligns with previous research [7, 13, 4]. Notably, the channel connections (TP8, T7), (FT8, T7), and (T8, FT7) reveal a strong correlation between the left and right hemispheres. The connection (FC6, C6) in head-2 shows the strongest activity, further indicating that local information in the central region may play a key role in emotion recognition.
For the SEED-IV dataset, the strongest inter-channel connections occur in the frontal, parietal, and occipital regions, consistent with existing findings [13]. In MVGT-G, the main channel connections are (O1, PO3), (P4, P2), (CB1, PO7), and (PO5, PO7). Additionally, the connection between AF3 and FP1 provides critical insights into emotion processing, in line with [13, 4].
Our model incorporates both intra-regional and inter-regional brain dynamics, highlighting the critical role of distributed network interactions in modeling emotional states, as discussed by [29].
IV Conclusions

### v5

@@ Abstract
1 Introduction
@@
Affective computing is commonly employed for the analysis of emotional states through Human-Computer Interaction (HCI) systems, which collect multimodal data from subjects, including voice signals, self-reports, body gestures and physiological signals. Among these modalities, physiological signals offer distinct advantages as they are directly derived from the subjects’ mental states, making it difficult for them to disguise or conceal their emotions. Scalp electroencephalography (EEG), a physiological signal usually used to analyze the cognitive functions of the human brain, is collected using noninvasive electrodes on the scalp [alarcao2017emotions]. Nowadays, due to its high temporal resolution, portability, and affordability, this method is widely employed to study brain changes in response to emotional stimuli [niemic2004studies].
@@
EEG signals contain rich emotional information in the temporal, frequency, and spatial domains. How to reasonably and effectively utilize information from each domain is a major challenge. Therefore, recent studies have attempted to combine multi-domain information to capture complex EEG features [MD-AGCN, EmoGT, li2022multi]. Although these works have achieved good performance, their limited use of domain-specific information has constrained model performance.
@@
The frequency domain features of EEG commonly include power spectral density (PSD), differential entropy (DE), differential asymmetry (DASM), and others. Among these, DE has been proven to be the simplest and most effective in EEG emotion recognition tasks [duan2013differential]. DE has the ability to distinguish EEG patterns between high-frequency and low-frequency energy [zheng2015investigating].
@@
Given the high temporal resolution of EEG signals, many studies use temporal models to extract sequential information. Recurrent neural networks (RNNs) have significantly improved emotion recognition performance by modeling temporal dependencies [R2G-STNN, BiDANN]. However, their serial structure poses challenges for parallelization. Convolutional neural networks (CNNs) have also been applied to extract temporal features, but their performance is restricted by limited receptive fields [lggNet, MD-AGCN]. To overcome these difficulties, [EmoGT] employs the attention mechanism to capture relationships between time points; however, this method may suffer performance degradation due to time-unaligned events at a single moment.
@@
The spatial domain of EEG is equally crucial for understanding emotional states, as emotions often involve distributed neural circuits rather than isolated regions [mauss2010measures]. Asymmetric activity between the left and right hemispheres can reflect changes in valence and arousal [schmidt2001frontal]. Message-passing-based graph neural network (MPGNN) methods can effectively capture structural relationships between EEG channels [rgnn, DGCNN, lggNet, MD-AGCN]. However, using MPGNN may pose potential risks of over-smoothing and over-squashing, which could hinder its ability to learn complex representations [ller2024attending]. Additionally, these methods often underutilize the geometric and anatomical structural relationships inherent in EEG spatial information.
Figure 1: Overall structure of MVGT. (a) represents the process of brain region encoding, centrality encoding and geometric structure encoding. (b) depicts the process of calculating inter-channel correlations based on the attention mechanism and geometric structure encoding. “Recycling” refers to the iterative refinement (see 2.4).
@@
To address these limitations and fully exploit features in the temporal, frequency, and spatial domains, we propose a multi-view graph transformer (MVGT) based on spatial relations for EEG emotion recognition. Graph transformers offer robust and flexible spatial perception while mitigating issues like over-smoothing and over-squashing [ller2024attending]. For the frequency domain, we use the embedding of DE as the feature representation. For the temporal domain, inspired by [itransformer], we expand the temporal receptive field by embedding entire continuous time segments as tokens and employ feed-forward neural networks to capture temporal patterns. In the spatial domain, we incorporate several simple yet effective spatial encodings, brain region encoding, centrality encoding and geometric structure encoding, into the model to help it capture effective spatial structures. These encodings allow MVGT to adaptively model correlations between EEG channels, delivering improved performance and deeper insights into emotional states.
2 Methods
2.1 Problem Definition
@@
2.2 Temporal Embedding
Temporal information is critical in time-series processing. [EmoGT] treated the features of different channels at the same time points as tokens and employed the attention mechanism to extract temporal correlations between them. However, due to anisotropic volume conduction characteristics [craik2019deep] in human brain tissues, there may be different physical meanings between channels at the same time step, resulting time-unaligned events that subsequently degrade performance. Other approaches [MD-AGCN, lggNet] utilized CNNs to extract temporal information along the time axis from continuous EEG segments, with the receptive field limited by the size of the convolution kernel. Inspired by [itransformer], we broaden the receptive field by embedding entire time segments as tokens rather than single time points. EEG signals are segmented using overlapping sliding windows of size $T$, and these segments are treated as tokens. After processing with sliding windows, we obtain $\tilde{\bm{X}}=\left(\tilde{\bm{X}}_{1},\tilde{\bm{X}}_{2},\cdots,\tilde{\bm{X}}_{S}\right)\in\mathbb{R}^{S\times n\times Tf}$, where $S$ denotes the number of continuous EEG segments, $n$ is the number of channels, and $f$ denotes the dimension of DE features. Refer to the Appendix 5.2 for more temporal embedding details.
According to the universal approximation theorem [hornik1991approximation], the feed-forward neural network (FFN), as the basic module of the Transformer [Transformer] encoder, can learn the intrinsic properties to describe a time series. By processing continuous time segments as input, the FFN extracts temporal features independently for each channel, providing more advantageous predictive representations compared to self-attention [itransformer].
2.3 Spatial Encoding
@@
2.3.1 Brain Region Encoding
Neuroscience research indicates that specific brain regions activate together during high-level cognitive processes [kober2008functional]. In EEG emotion recognition, incorporating relevant neuroscience findings can typically enhance recognition accuracy. For example, studies like [rgnn] and [BiHDM] utilized the asymmetry of neural activity between the left and right hemispheres to enhance recognition performance. [lggNet] divided EEG channels into different regions and combined intra-region convolution with inter-region convolution to capture spatial relationships. Building on [lggNet], we propose four learnable BRE schemes based on neuroscience research, ranging from coarse-grained to fine-grained encodings, namely: LOBE, FRONTAL, GENERAL, and HEMISPHERE (see Appendix 5.4 for design ideas). Coarse-grained region division can aggregate more intra-region information, while fine-grained division can capture more inter-region information. The four schemes mentioned above are shown in Fig. 2.
@@ 2.3.1 Brain Region Encoding
2.3.2 Structure Encoding
EEG channels are distributed in the 3D space, and the functional connectivity between them remains imprecisely defined. Therefore, inspired by [graphormer3D], we model the relationships between EEG channels as a fully connected directed graph to avoid making specific assumptions about functional connectivity. Firstly, let $\phi(i,j)$ represent the Euclidean distance between node $i$ and node $j$, and encode $\phi(i,j)$ using a set of Gaussian basis functions [graphormer3D, spinconv]. Let $\bm{b}_{k}\in\mathbb{R}^{n\times n}$ denote one of the Gaussian basis functions. The element $(i,j)$ of this function can be expressed as:
@@ 2.3.2 Structure Encoding
where $\mathbin{\|}$ denotes the concatenation operator. Secondly, the structure encoding is summed along the second dimension, followed by an edge projection through a linear layer to derive the CE (see (4)) that reflects each node’s relative importance. Nodes with higher cumulative weights are considered to have a greater level of involvement within the network. The encoding process for a node in a sample of the input data is as follows:
@@
Unlike previous symmetric adjacency matrices [EmoGT, rgnn], our directed spatial encoding matrix allows the model to learn distinct correlations for $\left(i,j\right)$ and $\left(j,i\right)$. Letting $l$ denote the model depth, $s$ denote the sample index and $m$ denote the index of multi-head attention, the multi-head attention module (MHA) can be represented as:
@@
2.4 Implementation Details of MVGT
In this section, we describe the overall architecture of the model, as illustrated in Fig. 1. For better optimization, we first apply GraphNorm [graphnorm] to normalize the input features to a range between 0 and 1. The encodings could be characterized as below:
@@ 2.4 Implementation Details of MVGT
We employ a Pre-LN structure, applying layer normalization (LN) before the MHA and the FFN. This design choice is supported by prior research [xiong2020on], which demonstrates that Pre-LN structures produce more stable gradients, facilitating faster and more reliable convergence. To further prevent overfitting, dropout is employed during training. The process is expressed mathematically as:
@@
Inspired by [graphormer3D, AlphaFold], we apply iterative refinement by recursively feeding the outputs of the model back into the same modules, denoted as “recycling” in Fig. 1. The iterative approach refines the model’s ability to discriminate encoded information and understand emotional patterns, thereby helping the model capture more effective details.
3 Experiments
3.1 Datasets and Settings
For our experiments, we select the SEED [zheng2015investigating] and SEED-IV [seediv] datasets to evaluate the effectiveness of our model. These datasets consist of EEG signals recorded from subjects watching emotion-eliciting videos. Following the settings of previous studies [zheng2015investigating, BiHDM, DGCNN, rgnn, MD-AGCN, EmoGT, seediv, li2022multi], we use pre-computed differential entropy (DE) features for the recognition task. Time segments are extracted using an overlapping sliding window of size $T$, consistent with [MD-AGCN], with $T$ set to 5. For the SEED dataset, we use the first 9 trials of each subject as the training set and the last 6 trials as the test set, as done in previous research. The DE features are computed using five frequency bands extracted from 1s nonoverlapping windows. The model performance is evaluated based on the average accuracy and standard deviation across all subjects over two sessions of EEG data. Similarly, for the SEED-IV dataset, we allocate the first 16 trials to training and the last 8 trials to testing. The DE features for SEED-IV are calculated using 4s windows. The performance of our model is assessed using data from all three sessions. Refer to Appendix 5.1 for DE feature extraction.
During experiments, the hidden dimension is set to 64 and the number of Gaussian basis functions is 32. The number of MHA layers is 4 and the number of attention heads is 2. We employ iterative refinement by recursively feeding the output back into the model for three times. We set the batch size to 32 and the learning rate within the range of 3e-5 to 3e-3. Cross-entropy is used as the loss function, and AdamW [adamw] is employed as the optimizer with a weight decay rate of 0.1.
Table 1: The classification accuracies (mean/std) on SEED and SEED-IV. MVGT-L, MVGT-G, MVGT-H, MVGT-F: MVGT using LOBE, GENERAL, HEMISPHERE and FRONTAL scheme. The best results are highlighted in bold, and the second-best results are underlined.
@@ 3.1 Datasets and Settings
DGCNN [DGCNN] 90.40/08.49 69.88/16.29
BiDANN [BiDANN] 92.38/07.04 70.29/12.63
BiHDM [BiHDM] 93.12/06.06 74.35/14.09
R2G-STNN [R2G-STNN] 93.34/05.96 -
RGNN [rgnn] 94.24/05.95 79.37/10.54
MD-AGCN [MD-AGCN] 94.81/04.52 87.63/05.77
EmoGT [EmoGT] 95.02/05.99 91.20/09.60
MV-SSTMA [li2022multi] 95.32/3.05 92.82/5.03
@@
3.2 Results Analysis
We compare the classification results on the SEED and SEED-IV datasets with recent state-of-the-art models, as shown in Table 1. Under identical experimental conditions, our proposed model demonstrated significantly better results than the baselines. On the SEED dataset, the FRONTAL scheme achieved the highest classification accuracy of 96.55%, representing a 1.23% improvement over the best baseline model. Similarly, on the SEED-IV dataset, the GENERAL scheme reached an accuracy of 94.03%, exceeding the best baseline by 1.21%. The differences in accuracy across region division schemes likely result from their data-dependent nature. This suggests that designing data-specific division schemes can enable the model to better utilize data characteristics, thereby enhancing recognition performance. Conversely, improper schemes may result in poor performance.
Fig. 3 shows the confusion matrices for MVGT-F on SEED and MVGT-G on SEED-IV.
@@ 3.2 Results Analysis
Table 2: Ablation study for the classification accuracies (mean/std) of MVGT-F on SEED and MVGT-G on SEED-IV. Symbol “✓” indicates the component is employed.
@@
For SEED, the model achieves 97.70% accuracy for positive emotions but performs less effectively on negative emotions (95.61%). Misclassification rates are low, with 0.76% of positive samples misclassified as negative and 1.08% of negative samples as positive, demonstrating effectiveness in distinguishing valence changes. For SEED-IV, the model achieves its best performance on neutral emotions (95.77%) but slightly lower accuracy on happy emotions (90.30%), likely due to the GENERAL scheme’s sensitivity to balanced emotions.
3.3 Ablation Study
To validate the effectiveness of our proposed method, we conduct ablation experiments on MVGT-F (based on SEED) and MVGT-G (based on SEED-IV). These experiments incrementally remove four key components: centrality encoding (CE), brain region encoding (BRE), geometric structure encoding (GSE), and the inverted temporal embedding (Inverted). GSE is incorporated into the attention mechanism as a bias (as detailed in (8)). The term “Inverted” refers to embedding continuous time segments as tokens (described in Section 2.2), whereas the default approach treats multi-channel information from a single time point as tokens.
@@ 3.3 Ablation Study
3.4 Visualization of Inter-channel Relations
@@
On the SEED dataset, the parameters suggest that emotional patterns involve activity across multiple brain regions. The parameters from head-1 in MVGT-F suggest that brain activity is primarily concentrated in the temporal and frontal regions on the lateral side of the brain, which aligns with previous research [zheng2015investigating, rgnn, EmoGT]. Notably, the channel connections (TP8, T7), (FT8, T7), and (T8, FT7) reveal a strong correlation between the left and right hemispheres. The connection (FC6, C6) in head-2 shows the strongest activity, further indicating that local information in the central region may play a key role in emotion recognition.
For the SEED-IV dataset, the strongest inter-channel connections occur in the frontal, parietal, and occipital regions, consistent with existing findings [rgnn]. In MVGT-G, the main channel connections are (O1, PO3), (P4, P2), (CB1, PO7), and (PO5, PO7). Additionally, the connection between AF3 and FP1 provides critical insights into emotion processing, in line with [rgnn, EmoGT].
Our model incorporates both intra-regional and inter-regional brain dynamics, highlighting the critical role of distributed network interactions in modeling emotional states, as discussed by [barrett2013large].
4 Conclusions